[Brahms-l] Shift report 20060617 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 17 2006 - 08:00:36 EDT
Shifter leader was cn
Shifters were nt


1.
We run
*FS 6 deg. 861 B (D1,D2,D3,D4) + MRS 90 deg. 700 B: 1 Full Stores 
*FS 6 deg. 861 A (D1,D2,D3,D4) + MRS 90 deg. 700 A: 2 Full Stores 
2.
The condition of
FS 6 deg. 861 B (D1,D2,D3,D4) + MRS 90 deg. 700 B: 1 Full Stores 
x 4 runs (~2 hrs).
The condition of
FS 6 deg. 861 A (D1,D2,D3,D4) + MRS 90 deg. 700 A: 2 Full Stores 
x 2 runs (~1 hr).
3.
The detectors are fine.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 17 08:01:19 2006

This archive was generated by hypermail 2.1.8 : Sat Jun 17 2006 - 08:01:28 EDT